OpenStudy (winsteria):

MEDAL AND FAN A piece of circular bread has a diameter of 11 in. What is the approximate area of this piece of bread? Use 3.14 to approximate pi and round your answer to the nearest hundredth.

OpenStudy (winsteria):

@StudyGurl14

OpenStudy (studygurl14):

Remember area of circle is \(r^2\pi\) r=radius diamter is twice the radius

OpenStudy (winsteria):

so what do I do to get the answer? @StudyGurl14

OpenStudy (studygurl14):

The diameter in this question 11. So what's the radius?

OpenStudy (winsteria):

5.5? @StudyGurl14

OpenStudy (studygurl14):

Correct. Now plug that into the formula given up there ^ and solve

OpenStudy (studygurl14):

remember to use 3.14 for pi

OpenStudy (winsteria):

what's the steps to solve it? @StudyGurl14

OpenStudy (studygurl14):

Plug in and solve. \(\Large r^2\pi\) r = 5.5 pi = 3.14 Thus, \(\Large (5.5)^2(3.14)\) You can do the rest.

OpenStudy (winsteria):

94.985? @StudyGurl14

OpenStudy (studygurl14):

yep.

OpenStudy (studygurl14):

round to the nearest hundredth though.

OpenStudy (winsteria):

90? @StudyGurl14

OpenStudy (studygurl14):

No.